# From affine algebraic racks to Leibniz algebras

### v1

@@ 8. Open questions
    Does the Leibniz algebra of a (not necessarily affine) algebraic rack have a natural Leibniz algebra structure? That is, does the functor ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})\to{\mathsf{Leib}_{k}}$ from Theorem 1.1 extend to a functor ${\mathsf{Rack}_{\ast}}({\mathsf{Sch}_{k}})\to{\mathsf{Leib}_{k}}$?
@@
    If $G$ is an algebraic group, then its Lie algebra $\mathrm{Lie}(G)$ is isomorphic to the space of left-invariant $k$-derivations $D\colon\mathcal{O}(G)\to\mathcal{O}(G)$, that is, $k$-derivations that satisfy the identity $\Delta\circ D=(\mathord{\operatorname{id}}\otimes D)\circ\Delta$. (See, for example, [milne]*Sec. 10e.) Do Leibniz algebras of algebraic racks have a similar interpretation?
@@
    Mirroring the representation theory of algebraic groups and comodules of commutative Hopf algebras (see, for example, [milne]*Sec. 4e), develop a theory of affine representations of affine algebraic racks $Q$ and $\mathcal{O}(Q)$-comodules. Also, study their relationships with the induced omni-representations (see [omni]) of Leibniz algebras. See Comment 7.3 for further discussion.
@@
    Building upon the previous question, introduce a cohomology theory for algebraic racks and their representations. Study its relationships with Leibniz algebra cohomology [lp], traditional rack and quandle cohomology [covez2], and (in the case of conjugation quandle schemes) the cohomology of algebraic groups and their representations (see [bendel]).

### v2

@@ 14. Open questions
    If $Q$ is any algebraic rack (not necessarily affine), then does $T_{e}Q$ have canonical Leibniz brackets? That is, do the functors ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})\to{\mathsf{Leib}_{k}}$ and ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})\to{\mathsf{RLeib}_{k}}$ from Theorem 1.1 extend to functors ${\mathsf{Rack}_{\ast}}({\mathsf{Sch}_{k}})\to{\mathsf{Leib}_{k}}$ and ${\mathsf{Rack}_{\ast}}({\mathsf{Sch}_{k}})\to{\mathsf{RLeib}_{k}}$?
@@
    If $G$ is an algebraic group, then its Lie algebra $\mathrm{Lie}(G)$ is isomorphic to the space of left-invariant $k$-derivations $D\colon\mathcal{O}(G)\to\mathcal{O}(G)$, that is, $k$-derivations that satisfy the identity $\Delta\circ D=(\mathord{\mathrm{id}}\otimes D)\circ\Delta$. (See, for example, [milne]*Sec. 10e.) Do Leibniz algebras of algebraic racks have a similar interpretation?
@@
    Mirroring the theory of representations of algebraic groups and comodules over commutative Hopf algebras (see, for example, [milne]*Sec. 4e), develop a theory of affine representations of affine algebraic racks and “comodules” over corack algebras. Also, study their relationships with the induced omni-representations (see [omni]) of Leibniz algebras. See Comment 13.3 for further discussion and references.
@@
    Building upon the previous question, develop a cohomology theory for algebraic racks and their representations. Study its relationships with Leibniz algebra cohomology (see [leibniz]), rack and quandle cohomology (see [book]), and (for conjugation quandle schemes) the cohomology of algebraic groups and their representations (see [milne]).